Al(OH)3(s) → Al3+(aq) + 3OH−(aq)

ΔfH∘/kJ⋅mol−1 −1284    −538.4 −230.02

S∘/J⋅mol−1⋅K−1 85.4 −325 -10.9

Part A Compute ΔrS (answer in J⋅mol−1⋅K−1)

Part B Compute ΔrG∘ at 298 K (answer in kJ⋅mol−1)

Part C Determine the equilibrium constant for the dissolution of Al(OH)3 at 298 K.

Part D Determine the Gibbs energy change for the dissolution of Al(OH)3 when the concentrations of Al3+ and OH− are both 2.60×10−7mol⋅L−1.(answer in kJ⋅mol−1)

Part E At what temperature will Al(OH)3(s) be at equilibrium with Al3+ and OH−, each at 2.60×10−7mol⋅L−1? (answer in K)
Answer #1

Al(OH)2 (A) = A13+ (aq) + 3 OH(aq) Aysº = Sproducts - Esreactants = (-325 + 3(-10.) - (85.4)] ank mol. K = -443 Jk mol-! B)[A13+][OH)3 Qc = = (2.60 x 10-7)(2.60 x107) = 4.57 X10-27 AG = Da + RT In Qc = . k 187.6 kJ mo I + 8314 J x 298 K x In 4.57X1